OpenStudy (bloomlocke367):

@rational Betty's Bite-Size Candies are packaged in bags. The number of candies per bag is normally distributed, with a mean of 50 and a standard deviation of 3. At a checkpoint, a sample of bags is checked, and 4 bags contain fewer than 47. How many bags were probably taken as samples?

OpenStudy (bloomlocke367):

What do I do here?

OpenStudy (bloomlocke367):

@rational

OpenStudy (rational):

We have standard deviation, \(\sigma = 3\) Notice that 47 is 1 \(\sigma \) less than the mean, 50.

OpenStudy (rational):

so the probability for one bag to have less than 47 candies would be : 13.6+2.1+0.1 = 15.8 %

OpenStudy (bloomlocke367):

okay..... I'm still confused, sorry

OpenStudy (rational):

15.8% = 4 can you guess what 100% would be ?

OpenStudy (rational):

its okay, you will understand these after doing 2-3 prblems

OpenStudy (rational):

0.158x = 4 solve x

OpenStudy (bloomlocke367):

I got 25.316... I did a proportion:\[\frac{ 15.8 }{ 4 }=\frac{ 100 }{ ? }\]

OpenStudy (rational):

yes so approximately 25 bags are taken as sample

OpenStudy (bloomlocke367):

oh, wow. okay

OpenStudy (rational):

dont watch it, its very advanced video

OpenStudy (rational):

il delete it

OpenStudy (bloomlocke367):

okay...